OpenStudy (anonymous):

Find the standard equation of the ellipse with the given characteristics and the center at the origin. Vertical major axis and passes through the point (0,6) & (3,0). Please HELP!! Will give Medal!!

OpenStudy (dumbcow):

 a = 6 (vertical y direction) b = 3 (horizontal x direction) \[\frac{x^2}{9} + \frac{y^2}{36} = 1\]